OpenStudy (anonymous):

what is the equation of a line that has a slope of -1 and passes through the point(-3,12)??

OpenStudy (mertsj):

Do you know point slope form of the equation of a line?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

dont i need two points??

OpenStudy (mertsj):

replace ysub1 with 12. replace m with -1. replace xsub1 with -3 and solve for y

OpenStudy (anonymous):

ok im a little lost now

OpenStudy (anonymous):

wont i still have x?

OpenStudy (mertsj):

\[y-12=-1(x- -3)\]

OpenStudy (mertsj):

\[y-12=-1(x+3)\]

OpenStudy (mertsj):

\[y-12=-x-3\]

OpenStudy (mertsj):

y=-x+9

OpenStudy (anonymous):

ohhh ok i understand now thanks again..

OpenStudy (mertsj):

And that is the equation of a line whose slope is -1 and passes through the point (-3,12).